OpenStudy (anonymous):

what is the value of cotx-sin(pi/2-x)cos(pi/2-x) (i)cotx cos^2x (ii)cot^2x (iii)cos^2 x (iv)tan^2x

OpenStudy (vishweshshrimali5):

\[\large{\sin (\cfrac{\pi}{2} - x) = \cos x}\] \[\large{\cos (\cfrac{\pi}{2} - x) = \sin x}\] Using this can you simplify the LHS ?

OpenStudy (anonymous):

is that a identity

OpenStudy (anonymous):

or how can we prove it

OpenStudy (vishweshshrimali5):

Actually the (pi/2 - x) things are no such identities ..... but it is useful to remember them.